# Bench-2-CoP: Can We Trust Benchmarking for EU AI Compliance?

### v1

@@ Abstract
The rapid advancement of General Purpose AI (GPAI) models necessitates robust evaluation frameworks, especially with emerging regulations like the EU AI Act and its associated Code of Practice (CoP). Current AI evaluation practices depend heavily on established benchmarks, but these tools were not designed to measure the systemic risks that are the focus of the new regulatory landscape. This research addresses the urgent need to quantify this "benchmark-regulation gap." We introduce Bench-2-CoP, a novel, systematic framework that uses validated LLM-as-judge analysis to map the coverage of 194,955 questions from widely-used benchmarks against the EU AI Act’s taxonomy of model capabilities and propensities. Our findings reveal a profound misalignment: the evaluation ecosystem is overwhelmingly focused on a narrow set of behavioral propensities—such as ’Tendency to hallucinate’ (53.7% of the corpus) and ’Discriminatory bias’ (28.9%)—while critical functional capabilities are dangerously neglected. Crucially, capabilities central to loss-of-control scenarios, including evading human oversight, self-replication, and autonomous AI development, receive zero coverage in the entire benchmark corpus. This translates to a near-total evaluation gap for systemic risks like ’Loss of Control’ (0.4% coverage) and ’Cyber Offence’ (0.8% coverage). This study provides the first comprehensive, quantitative analysis of this gap, offering critical insights for policymakers to refine the CoP and for developers to build the next generation of evaluation tools, ultimately fostering safer and more compliant AI.
@@ 4 Results
Our analysis of 194,955 benchmark questions across six major evaluation suites reveals a stark and consequential misalignment between current AI evaluation practices and the systemic risk framework of the EU AI Act. The results demonstrate that while benchmarks provide extensive coverage for a few specific behavioral propensities, they leave critical functional capabilities almost entirely unevaluated. This creates a landscape of deep but narrow evaluation, ill-suited for comprehensive regulatory assessment and proactive risk management. A crucial methodological consideration underpins our analysis: the vast difference in scale between benchmarks, from MMLU’s over 100,000 questions to TruthfulQA’s 790. We report absolute question counts to convey the sheer volume and focus of the evaluation ecosystem. However, our analysis also emphasizes the relative, percentage-based focus within each benchmark, as this reveals the design priorities and conceptual commitments of each evaluation suite, preventing the scale of larger benchmarks from obscuring the contributions of smaller, more specialized ones.
@@ 4.1 Overall Coverage of the Benchmark Ecosystem
An aggregate view of the entire corpus immediately exposes a fundamental imbalance: the evaluation ecosystem is overwhelmingly focused on assessing model propensities (behavioral tendencies) rather than their underlying capabilities (functional abilities). As shown in Table 6, three propensities alone—P3 (Tendency to hallucinate), P4 (Discriminatory bias), and P5 (Lack of performance reliability)—are addressed in questions that collectively make up the vast majority of our corpus. This finding is significant, as it suggests that the industry’s evaluation efforts are primarily directed at managing the observable, and often publicly criticized, "symptoms" of model failure (like making things up or showing bias) rather than assessing the underlying functionalities that could lead to more novel or systemic harms. In stark contrast, most capabilities identified as critical by the CoP, particularly those related to autonomy and self-modification, receive minimal to zero coverage.
@@
This propensity-heavy distribution likely reflects the historical evolution of AI safety concerns, which focused on observable failures like bias and misinformation long before the rise of highly autonomous models. Furthermore, propensities like hallucination are methodologically easier to evaluate with static question-answering formats than complex capabilities like long-horizon planning or self-improvement, which require more interactive and dynamic testing environments. The result, however, is an evaluation landscape that is reactive to past failures rather than proactive about future risks. This reactive posture is a critical shortcoming in the context of the forward-looking EU AI Act, which is explicitly designed to anticipate and mitigate risks from future, more powerful systems.
@@
Category Type Total Questions % of Corpus
@@
Tendency to hallucinate (P3) Propensity 104 640 53.7%
Discriminatory bias (P4) Propensity 56 249 28.9%
Lack of performance reliability (P5) Propensity 36 258 18.6%
@@
Process multiple modalities (C11) Capability 9 502 4.9%
Tendency to deploy harmfully (P2) Propensity 1 857 1.0%
Lawlessness (P6) Propensity 1 353 0.7%
@@
Long-horizon planning (C6) Capability 851 0.4%
Misalignment with human intent (P1) Propensity 813 0.4%
Manipulate, persuade, deceive (C3) Capability 642 0.3%
Control physical systems (C13) Capability 277 0.1%
Use tools/computer use (C12) Capability 191 0.1%
Adaptive learning (C5) Capability 169 0.1%
@@
CBRN capabilities (C2) Capability 17 0.01%
Goal-pursuing/power-seeking (P7) Propensity 15 0.01%
Self-reasoning (C7) Capability 11 0.006%
Autonomy (C4) Capability 5 0.003%
@@ 5.2 Methodological Considerations and Study Limitations
It is also important to state that our analysis, while deep, is not exhaustive of the entire evaluation landscape. Our corpus was constructed from a curated selection of six influential benchmarks, chosen for their widespread use by major developers and their relevance to systemic risk. This was a deliberate scoping decision to allow for a detailed and rigorous analysis of the most prominent evaluation tools. However, this means our study does not capture the full diversity of all publicly available benchmarks. A clear and valuable direction for future work is to apply our Bench-2-CoP framework to a broader set of existing benchmarks to create an even more comprehensive map of the evaluation ecosystem.
Finally, our analysis is necessarily focused on publicly documented benchmarks, excluding any proprietary, internal evaluations that organizations may conduct. While this limitation means we cannot capture the complete landscape of AI safety evaluation, the benchmarks we studied represent the primary public evidence used for safety assurance and are the tools most commonly cited in technical reports and regulatory submissions. Their limitations are therefore highly consequential for public accountability, independent auditing, and regulatory oversight.
5.3 Lessons Learned
@@ 5.4 Future Work: Building the Next Generation of Regulatory-Aligned Benchmarks
The gaps this study has quantified point directly to a clear and urgent agenda for future work. A foundational next step, as noted in our limitations, is to expand the application of the Bench-2-CoP framework to a wider array of public benchmarks, creating an ever-more-complete picture of the evaluation landscape. Beyond this, our findings strongly advocate for two complementary paths for new benchmark creation.

### v2

@@ Abstract
The rapid advancement of General Purpose AI (GPAI) models necessitates robust evaluation frameworks, especially with emerging regulations like the EU AI Act and its associated Code of Practice (CoP). Current AI evaluation practices depend heavily on established benchmarks, but these tools were not designed to measure the systemic risks that are the focus of the new regulatory landscape. This research addresses the urgent need to quantify this "benchmark-regulation gap." We introduce Bench-2-CoP, a novel, systematic framework that uses validated LLM-as-judge analysis to map the coverage of 194,955 questions from widely-used benchmarks against the EU AI Act’s taxonomy of model capabilities and propensities. Our findings reveal a profound misalignment: the evaluation ecosystem dedicates the vast majority of its focus to a narrow set of behavioral propensities. On average, benchmarks devote 61.6% of their regulatory-relevant questions to ’Tendency to hallucinate’ and 31.2% to ’Lack of performance reliability’, while critical functional capabilities are dangerously neglected. Crucially, capabilities central to loss-of-control scenarios, including evading human oversight, self-replication, and autonomous AI development, receive zero coverage in the entire benchmark corpus. This study provides the first comprehensive, quantitative analysis of this gap, demonstrating that current public benchmarks are insufficient, on their own, for providing the evidence of comprehensive risk assessment required for regulatory compliance and offering critical insights for the development of next-generation evaluation tools.
@@ 4 Results
Our analysis of 194,955 benchmark questions across six major evaluation suites reveals a stark and consequential misalignment between current AI evaluation practices and the systemic risk framework of the EU AI Act. The results demonstrate that while benchmarks provide extensive coverage for a few specific behavioral propensities, they leave critical functional capabilities almost entirely unevaluated. This creates a landscape of deep but narrow evaluation, ill-suited for comprehensive regulatory assessment. To present the most accurate picture, our analysis considers two perspectives: the raw volume of questions, which shows the absolute focus of the ecosystem, and a normalized view that averages each category’s percentage contribution within each benchmark. This second view provides a fairer comparison of design priorities, preventing the sheer scale of larger benchmarks like MMLU from overshadowing the intent of smaller, more specialized ones.
@@ 4.1 Overall Coverage of the Benchmark Ecosystem
An aggregate view of the entire corpus immediately exposes a fundamental imbalance: the evaluation ecosystem is overwhelmingly focused on assessing model propensities rather than their underlying capabilities. As shown in Table 6, this holds true even when normalizing for benchmark size. The data reveals that, on average, benchmarks dedicate the lion’s share of their regulatory-relevant questions to a few key propensities. ’Tendency to hallucinate’ (P3) remains dominant, with an average of 61.6% of each benchmark’s focus.
This normalized view, however, tells a more nuanced story than raw counts alone. For example, ’Discriminatory bias’ (P4), while second in raw numbers, is third in normalized priority (17.2%), reflecting the fact that its high absolute count is driven almost entirely by one specialized benchmark (BBQ). Conversely, ’Tendency to deploy harmfully’ (P2) has a very low raw percentage (0.95%) but a much higher average normalized percentage (12.35%). This indicates that while few benchmarks test for it, those that do (primarily HLE) consider it a very high priority. This highlights a key finding: the ecosystem’s evaluation of certain critical risks is highly fragmented and dependent on a few niche, forward-looking benchmarks rather than being an integrated, widespread practice.
@@
This propensity-heavy distribution likely reflects the historical evolution of AI safety concerns, which focused on observable failures like bias and misinformation long before the rise of highly autonomous models. The result, however, is an evaluation landscape that is reactive to past failures rather than proactive about future risks—a critical shortcoming in the context of the forward-looking EU AI Act.
@@
Category Type Total Questions Avg. Normalized %
@@
Tendency to hallucinate (P3) Propensity 104 640 61.62%
Discriminatory bias (P4) Propensity 56 249 17.20%
Lack of performance reliability (P5) Propensity 36 258 31.19%
@@
Process multiple modalities (C11) Capability 9 502 7.86%
Tendency to deploy harmfully (P2) Propensity 1 857 12.35%
Lawlessness (P6) Propensity 1 353 0.78%
@@
Long-horizon planning (C6) Capability 851 3.04%
Misalignment with human intent (P1) Propensity 813 0.16%
Manipulate, persuade, deceive (C3) Capability 642 0.93%
Control physical systems (C13) Capability 277 0.15%
Use tools/computer use (C12) Capability 191 0.14%
Adaptive learning (C5) Capability 169 0.13%
@@
CBRN capabilities (C2) Capability 17 0.004%
Goal-pursuing/power-seeking (P7) Propensity 15 0.10%
Self-reasoning (C7) Capability 11 0.23%
Autonomy (C4) Capability 5 0.002%
@@ 5.2 Methodological Considerations and Study Limitations
It is also important to state that our analysis, while deep, is not exhaustive of the entire evaluation landscape. Our corpus was constructed from a curated selection of six influential benchmarks. This was a deliberate scoping decision to focus on the benchmarks most commonly disclosed by major commercial AI developers in their public-facing model cards and technical reports. The purpose was not to claim that no other safety benchmarks exist, but to demonstrate the gap between what is required for compliance and what is being publicly used as evidence. While we are aware that companies engage in more specialized, often confidential, external evaluations (e.g., with organizations like METR or Apollo), the very fact that these practices are not part of the standardized, public accountability framework reinforces our central point about the compliance gap.
Finally, our approach to mapping capabilities and propensities to systemic risks was a necessary methodological choice. Since the Code of Practice does not provide an explicit mapping, our expert-led construction of "risk pathways" was a required step to operationalize the regulation for quantitative analysis. Our key finding remains that commercial AI evaluation practices, as currently disclosed, leave massive blind spots for CoP-critical capabilities.
5.3 Recommendations for Key Stakeholders
@@ 5.4 Future Work: Building the Next Generation of Regulatory-Aligned Benchmarks
The gaps this study has quantified point directly to a clear and urgent agenda for future work. A foundational next step is to expand the application of the Bench-2-CoP framework to a wider array of public benchmarks, creating an ever-more-complete picture of the evaluation landscape. Beyond this, our findings strongly advocate for two complementary paths for new benchmark creation.